# ToDo (#1): #2

### v3

@@ 2.3.1 Technical Design
The technical design was presented in Refs. [ratzingerTechnicalDesignReport2020, winklehner_high-current_2021, holtermann_technical_2021] and the optimization of the RF loop coupler in Ref. [sangroula_design_2021]. The thermal properties and cooling considerations were presented in Ref. [koser_thermal_2021]. An optimization scheme using machine learning was presented in Ref. [koser_input_2022].
Figure 2.12: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz. From Ref. [holtermann_technical_2021].
In terms of mechanical stability, our analyses show two quarter wave modes which are associated with the choice of a split-coaxial resonator. The addition of bridges between opposite vanes (see Fig. 2.12) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
@@ 2.3.2 Beam Dynamics
Beam dynamics studies were performed of the full 3D particle distribution resulting from the LEBT simulations going through the RFQ, including the effects of the entrance gap and asymmetry (due to the split coaxial design) and the exit gap and asymmetry. This was performed by the subcontracted company Bevatech GmbH in Germany, who is currently also manufacturing the RFQ. The well-established tracking code TraceWin was used [tracewin]. Fine tuning of the phase of the exit gap (i.e. tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Tab. 2.4.
Table 2.4: RFQ output beam parameters.
@@
$\varepsilon_{x}$ (1-RMS, norm.) 0.54 mm-mrad
$\varepsilon_{y}$ (1-RMS, norm.) 0.52 mm-mrad
$\varepsilon_{z}$ (1-RMS, norm.) 0.114 MeV-deg
Phase width (1-RMS) 40
@@ 2.3.3 Risks and Mitigation
Chapter 3 Design of the Cyclotron
Figure 3.1: Isometric view of the cyclotron.
In this section, we describe the details of the cyclotron at the level of a preliminary design report. This section includes details of the design of the cyclotron, the spiral inflector that transmits ions from the RFQ into the cyclotron, and the extraction system.
This section also characterizes the latest in simulo performance of these systems.
@@ 3.1 Magnet Yoke
The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron (seen in Fig 3.2), which is slightly larger, at 7 meters outer diameter.
Figure 3.2: C400 on the milling machine, C230 in front of it. Courtesy of IBA.
@@ 3.1.1 Design and Fabrication of the Cyclotron Magnet
As with the C400, the IsoDAR cyclotron is divided into smaller parts to facilitate transport to the site. The magnet is split radially and will be assembled on site. The vacuum chamber is split into 3 self-centering parts along its height. The vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers where 2 radial O-rings are used. An exploded visualization of the upper and lower parts is shown in Figs. 3.4 and 3.4. The material used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
Figure 3.3: Cyclotron upper steel exploded.
Figure 3.4: Cyclotron lower steel exploded.
@@
An cut-section view of the cyclotron is shown in Fig. 3.5. The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock). The poles are split into several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance. A detailed view of the cyclotron components is shown in Fig. 3.6 and additional details about the yoke parts can be found in Table 3.1
Figure 3.5: Cyclotron cut view
Figure 3.6: Cyclotron detailed cut view
Name Mass (Metric Ton) Dimensions (mm)
Upper and lower external disk 30 Ø4700*250
Upper and lower internal disk 36 Ø4620*463
Upper and lower flux return (1 and 2) 36 3100*1950*1015
Upper and lower median plane flux return (1 and 2) 29 3100*1950*705
Upper and lower pole 7 2013*1290*600
Table 3.1: Yoke part properties
3.2 Cyclotron magnetic performance
@@
In this section we describe the latest modifications applied to the cyclotron regarding its magnetic properties. We will compare the results with the model from the Conceptual Design Report (CDR) from [abs_isodarkamland_2015] to ensure its magnetic performance remains within specifications.
@@
With the latest design of the vacuum chamber, see section 3.5, the pole radius was slightly reduced. To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius required small modifications, see left image in Fig. 3.7. First, the curvature was modified to follow the curvature of the last orbit. This enhances the total flux in the region where needed. Second, the top layer of the pole at its outer radius will be made of permendur. This alloy has better magnetic properties at high field strengths. Using permendur will enhance the magnetic flux. Finally, the pole height is increased by a slanted ridge at its extremity. This reduction of the pole gap should have no effect on the transmission and can be tuned to improve the dB/dr for the last orbits.
@@
Figure 3.7: Left: Detailed view of the outer radius of the pole; only half of one pole is shown. In red is the coil, in light green is the pole made of SNVA6342 steel. The turquoise part is a piece 30 mm in height and 200 mm in width made of permendur. In blue we see the pole edge made of SNVA6342 steel, which can be shimmed to isochronize the magnetic field. Right: BH curves for the two steel types used: Permendur and SNVA6342. As reference, the BH curve used for the cyclotron in the CDR is also shown.
@@
The steel used for the cyclotron yoke and poles is the SNVA6342, also used, for example, in the C400 cyclotron (see Ref. [bib:c400_iron]). Its BH-curve is shown in Fig.3.7. Here, we also see the better performance of the permendur alloy.
@@ 3.2.1 Isochronization
The edge of the pole can be shimmed to isochronize the magnetic field map. In Fig. 3.8 the latest results are given. The left image shows the shimming applied, the right image shows the offset in magnetic field, as required for a given RF-frequency. For comparison, the offset (expressed as dB in Gauss) for the model of the CDR is also shown.
Figure 3.8: Left: Pole edge of isochronized model. Right: The dB [Gauss] as a function of radius [meters], for the CDR model (in black) and the newest model (in red). Note that a positive dB means there is too little field at specific radius, a negative dB implies there is too much field. The RF-frequency for the CDR model is at 32.77 MHz, for the newest model it is 32.82 MHz.
The plot shows that the magnetic field is stable and strong enough up to a radius of 2 meters, which is the equivalent of 121 MeV. In Fig 3.9 the phase slip can be seen to stay within the ±$\pm 10^{\circ}$. The horizontal and vertical tunes are also given in the same figure. Here we see that there is a crossing of the tunes at high radius, invoking resonances which could impact the extraction.
Figure 3.9: Left: phase slip as a function of radius. Right: vertical and horizontal tunes as a function of energy (dashed and solid lines respectively). Shown are the results for the CDR model (in black) and the newest model (in red).
@@ 3.3 Interface with the RFQ
Figure 3.10: RFQ overview.
The RFQ will sit on top of the cyclotron along its central axis, as shown in Fig 3.10. This places the RFQ outside of the cyclotron’s vacuum. Therefore, a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum. This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.12 ). The O-rings willhave a cross sectional diameter of 8mm and an inner diameter of 305mm.
In order to allow for adjustment, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ). The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the Yoke parts).
Figure 3.11: RFQ zoom.
Figure 3.12: RFQ cut section 1.
As shown in Fig. 3.13, vacuum sealing will occur at the base of the RFQ using a thick O-ring (diameter 10 mm) which will self-seal under the RFQ’s weight ($\approx 3000$ N). The dimensions of the vacuum sleeve are listed in Table 3.2.
In order to allow fine-tuning the cyclotron alignment, a set of adjustment feet, shown in Fig. 3.11, are fixed to the cyclotron yoke. By using four groups of set screws and one vertical adjustment nut, the following alignment modifications can be made.
  • Rotation: ±$\pm 5$
  • Tilt: ±$\pm 5$ mrad
  • Lateral movement: ±$\pm 1$ mm
@@
Figure 3.13: RFQ cut section 2.
Table 3.2: Vacuum Sleeve Dimensions.
@@ 3.4 Central Region
Figure 3.14: CAD rendering of the spiral inflector and particle trajectories as implemented previously for early tests at Best Cyclotron Systems, Inc. [alonso_isodar_2015]. The positive and negative electrostatic deflectors are shown in blue and green respectively. The cyclotron magnetic field in this image is directed vertically upwards. Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes. The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron. From Ref. [winklehner_realistic_2017].
The injection system design is a crucial component that allows us to achieve a high beam current. Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms. Higher energy injection mitigates space charge and reduces geometrical (non-normalized) emittance. An energy of 35 keV/amu was selected. This in turn determines the target output energy of the RFQ. As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane. We have investigated several refinements to a traditional spiral inflector which will be discussed subsequently, but the concept, as implemented successfully in a previous iteration of this work and tested at Best Cyclotron Systems, Inc., is visualized in Figure 3.14 using trajectories calculated with OPAL [adelmann_opal_2019].
@@
The central region and the spiral inflector must be designed carefully and together because they have strong interplay. This procedure has already undergone one past iteration in the summers of 2013 and 2014 at Best Cyclotron Systems Inc. In the following sections, we will discuss improvements to this design which will be incorporated into the preliminary design for the full IsoDAR machine.
@@ 3.4.1 Spiral Inflector Overview & Design Considerations
The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.14). The electrostatic potential between these electrodes is able to bend the beam 90 from the axial line to the median plane of the cyclotron. The helical trajectory of the beam is determined both by the shape of the electrodes, the electrode potentials, and by the magnetic field of the cyclotron as the beam is bent into the median plane. In order to isolate the inflector region from external fields as much as possible, grounded entrance and exit apertures and a grounded copper housing are also included.
@@ 3.4.2 Designing the Inflector
An initial estimate based on an analytical solution for a central track serves as a numerical starting point, after which the mesh generating software gmsh [GMSHarticle] is used to produce surface representations which inform subsequent electrostatic analysis. Python’s BEMPP API [bempp] is then used to perform boundary analysis.
In short, this procedure enables a geometric model of our inflector to yield a full 3D electrostatic. Particles can then be tracked with high precision using the Boris method [boris].
A major strength of this procedure and choice of software is speed, with a full simulation iteration taking only $\mathcal{O}$(hours). Prior work [weigelspiral] comparing this process to commercial software such as COMSOL demonstrates that the differences in field calculations are negligible.
@@ 3.4.3 Geometric Refinements
The addition of intermittant beam focusing ensures that the $\mathrm{H}_{2}^{+}$ entering the spiral inflector are as near to the central trajectory as possible. This is accomplished by the addition of electrostatic quadrupoles in the gap between the RFQ and inflector entrance (see also Fig. 1.7). Using the output of the RFQ simulation [rfqsim] as input to the quadrupoles, geometry and potentials were optimized such that the tightest possible input beam is produced.
The remaining three modifications: V-shape, plate angling, and the terminal wedge were motivated by work done by Barnard et al. [barnard_inflector]. These all have the effect of introducing quadrupole moments to the field of the inflector itself, which can provide mid-track focusing to minimize beam dispersion in flight.
The V-shape introduces a concave v-shape to the anode and a convex v-shape to the cathode when viewed along the long axis of the electrode (Fig.3.15(a)). Plate angling reorients the plate angle relativce to each other along the course of the inflector. In other words the plates are parallel only in the center with a positive angle between the plates initially linearly transitioning to an equal and opposite negative angle by the end (Fig.3.15(b)). The terminal gamma wedge cut modifies the end of the anode into a triangular rather than flat profile at the end and shortens the cathode. This assists with the fringe fields at the end which can produce vertical spreading. (Fig.3.15(c)).
@@
Figure 3.15: Illustrations of different focusing concepts. From Ref. [barnard_inflector].
@@
As we are at the edge of practical optimization using Monte Carlo, this leads us to consider other potential alternatives. Surrogate models using a fully connected neural network capable of learning and interpolating underlying features have been previously explored[EdelenMachingLearning] [KoserMachineLearning] [villarrealMachineLearning]. Initial iterations have proved promising and have informed the parameters for our current best inflector performance. As this work matures, it will be detailed in upcoming publications.

### v4

@@ 4.3.1 Technical Design
The technical design was presented in Refs. [45, 46, 47] and the optimization of the RF loop coupler in Ref. [48]. The thermal properties and cooling considerations were presented in Ref. [49]. An optimization scheme using machine learning was presented in Ref. [50].
Figure 20: Modes of the mechanical deformation of the RFQ vanes calculated with COMSOL. The frequencies are 18 Hz and 25 Hz. From Ref. [47].
In terms of mechanical stability, our analyses show two quarter wave modes which are associated with the choice of a split-coaxial resonator. The addition of bridges between opposite vanes (see Fig. 20) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
@@ 4.3.2 Beam Dynamics
Beam dynamics studies were performed of the full 3D particle distribution resulting from the LEBT simulations going through the RFQ, including the effects of the entrance gap and asymmetry (due to the split coaxial design) and the exit gap and asymmetry. This was performed by the subcontracted company Bevatech GmbH in Germany, who is currently also manufacturing the RFQ. The well-established tracking code TraceWin was used [51]. Fine tuning of the phase of the exit gap (i.e. tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Table 5.
Table 5: RFQ output beam parameters.
@@
$\epsilon_{x}$ (RMS, norm.) 0.54 mm-mrad
$\epsilon_{y}$ (RMS, norm.) 0.52 mm-mrad
$\epsilon_{z}$ (RMS, norm.) 0.114 MeV-deg
Phase width (RMS) 40
@@ 4.3.3 Risks and Mitigation
5 Design of the Cyclotron
Figure 21: Isometric view of the cyclotron.
In this section, we describe the details of the cyclotron at the level of a preliminary design report. This section includes details of the design of the cyclotron, the spiral inflector that transmits ions from the RFQ into the cyclotron, and the extraction system. This section also characterizes the latest simulated performance of these systems.
@@ 5.1 Magnet Yoke
The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron , which is slightly larger, at 7 m outer diameter [52].
@@ 5.1.1 Design and Fabrication of the Cyclotron Magnet
As with the C400, the IsoDAR cyclotron will be divided into smaller parts to facilitate transport to the site. The magnet is split radially and will be assembled on site. The vacuum chamber is split into 3 self-centering parts along its height. The vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers where 2 radial O-rings are used. An exploded visualization of the upper and lower parts is shown in Figs. 23 and 23. The material used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
Figure 22: HCHC-60 cyclotron upper steel exploded.
Figure 23: HCHC-60 cyclotron lower steel exploded.
@@
A cut-section view of the cyclotron is shown in Fig. 24. The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock). The poles are split into several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance. A detailed view of the cyclotron components is shown in Fig. 25 and additional details about the yoke parts can be found in Table 6
Figure 24: Cyclotron cut view
Figure 25: HCHC-60 cyclotron, detailed cut view calling out the major components of the steel, vacuum chamber, and coils.
Table 6: Steel yoke, individual parts masses and dimensions.
Name Mass (t) Dimensions (mm)
Upper and lower external disk 30 $4700\times 250$
Upper and lower internal disk 36 $4620\times 463$
Upper and lower flux return (1 and 2) 36 $3100\times 1950\times 1015$
Upper and lower median plane flux return (1 and 2) 29 $3100\times 1950\times 705$
Upper and lower pole 7 $2013\times 1290\times 600$
5.2 Simulated cyclotron magnetic performance
@@
In this section we describe the latest modifications applied to the cyclotron design regarding its magnetic properties. We will compare these results with the model from the Conceptual Design Report (CDR) from [3] to ensure its theoretical magnetic performance remains within specifications.
@@
With the latest design of the vacuum chamber, see Section 5.6, the pole radius was slightly reduced. To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius required small modifications, see left image in Fig. 26. First, the curvature was modified to follow the curvature of the last orbit. This enhances the total flux in the region where needed. Second, the top layer of the pole at its outer radius will be made of permendur. This alloy has better magnetic properties at high field strengths. Using permendur will enhance the magnetic flux. Finally, the pole height is increased by a slanted ridge at its extremity. This reduction of the pole gap should have no effect on the transmission and can be tuned to improve the dB/dr for the last orbits.
@@
Figure 26: Left: Detailed view of the outer radius of the pole; only half of one pole is shown. In red is the coil, in light green is the pole made of SNVA6342 steel. The turquoise part is a piece 30 mm in height and 200 mm in width made of permendur. In blue we see the pole edge made of SNVA6342 steel, which can be shimmed to isochronize the magnetic field. Right: BH curves for the two steel types used: Permendur and SNVA6342. As reference, the BH curve used for the cyclotron in the CDR is also shown.
@@
The steel used for the cyclotron yoke and poles is the SNVA6342, also used, for example, in the C400 cyclotron (see Ref. [53]). Its BH-curve is shown in Fig.26. Here, we also see the better performance of the permendur alloy.
@@ 5.2.1 Isochronization
The edge of the pole can be shimmed to isochronize the magnetic field map. In Fig. 27, the latest results are given. The left image shows the shimming applied, the right image shows the offset in magnetic field, as required for a given RF frequency. For comparison, the offset (expressed as dB in Gauss) for the model of the CDR is also shown.
[…]
The plot shows that the magnetic field is stable and strong enough up to a radius of 2 m, which is the equivalent of 121 MeV. In Fig. 28, the phase slip can be seen to stay within the $\pm 10^{\circ}$. The horizontal and vertical tunes are also given in the same figure. Here we see that there is a crossing of the tunes at high radius, invoking resonances which could impact the extraction.
Figure 28: Left: phase slip as a function of radius. Right: vertical and horizontal tunes as a function of energy (dashed and solid lines respectively). Shown are the results for the CDR model (in black) and the newest model (in red).
@@ 5.3 Interface with the RFQ
The RFQ will sit on top of the cyclotron along its central axis, as shown in Fig. 21. This places the RFQ outside of the cyclotron’s vacuum. Therefore, a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum. This sleeve will be fixed on top of the upper external disk, and the vacuum sealing will be done radially on the upper internal disk (see Fig. 29 ). The O-rings will have a cross-sectional diameter of 8 mm and an inner diameter of 305 mm.
To allow for adjustment, the RFQ sleeve permits a rotation of 5 mrad around the lower O-ring contact patch (as discussed with the RFQ manufacturer). The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the yoke parts).
Figure 29: CAD rendering of the RFQ inside the cyclotron yoke. The vacuum sleeve is called out in magenta.
Figure 30: CAD rendering of the RFQ inside the cyclotron yoke. The vacuum sleeve is called out in magenta and the O-ring forming the vacuum interface between RFQ and sleeve is zoomed in.
As shown in Fig. 30, vacuum sealing will occur at the base of the RFQ using a thick O-ring (diameter 10 mm) which will self-seal under the RFQ’s weight ($\approx 3000$ N). The dimensions of the vacuum sleeve are listed in Table 7.
To allow fine-tuning the cyclotron alignment, a set of adjustment feet, shown in Fig. 31, are fixed to the cyclotron yoke. By using four groups of set screws and one vertical adjustment nut, the following alignment modifications can be made.
  • Rotation: $\pm 5$
  • Tilt: $\pm 5$ mrad
  • Lateral movement: $\pm 1$ $m$
@@
Figure 31: CAD rendering of the top of the RFQ above the cyclotron top surface. The adjustment feet are called out in yellow.
Table 7: Vacuum Sleeve Dimensions.
@@ 5.4 Central Region
Figure 32: CAD rendering of the spiral inflector and particle trajectories as implemented previously for early tests at Best Cyclotron Systems, Inc. [39]. The positive and negative electrostatic deflectors are shown in blue and green respectively. The cyclotron magnetic field in this image is directed vertically upwards. Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes. The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron. From Ref. [54].
The injection system design is a crucial component that allows us to achieve a high beam current. Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms. Higher energy injection mitigates space charge and reduces geometrical (non-normalized) emittance. An energy of 35 keV/amu was selected. This in turn determines the target output energy of the RFQ. As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane. We have investigated several refinements to a traditional spiral inflector which will be discussed subsequently, but the concept, as implemented successfully in a previous iteration of this work and tested at Best Cyclotron Systems, Inc. (BCSI), is visualized in Fig. 32 using trajectories calculated with OPAL [55].
@@
The central region and the spiral inflector must be designed carefully and together because they have strong interplay. This procedure has already undergone one past iteration in the summers of 2013 and 2014 at BCSI. In the following sections, we will discuss improvements to this design which will be incorporated into the preliminary design for the full IsoDAR machine.
@@ 5.4.1 Spiral Inflector Overview & Design Considerations
The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Fig. 32). The electrostatic potential between these electrodes is able to bend the beam 90 from the axial line to the median plane of the cyclotron. The helical trajectory of the beam is determined both by the shape of the electrodes, the electrode potentials, and by the magnetic field of the cyclotron as the beam is bent into the median plane. In order to isolate the inflector region from external fields as much as possible, grounded entrance and exit apertures and a grounded copper housing are also included.
@@ 5.4.2 Designing the Inflector
An initial estimate based on an analytical solution for a central track serves as a numerical starting point, after which the mesh generating software Gmsh [56] is used to produce surface representations which inform subsequent electrostatic analysis. Python’s BEMPP API [57] is then used to perform boundary analysis.
In short, this procedure enables a geometric model of our inflector to yield a full 3D electrostatic. Particles can then be tracked with high precision using the Boris method [58].
A major strength of this procedure and choice of software is speed, with a full simulation iteration taking only $\mathcal{O}$(hours). Prior work [59] comparing this process to commercial software such as COMSOL demonstrates that the differences in field calculations are negligible.
@@ 5.4.3 Geometric Refinements
The addition of intermittant beam focusing ensures that the $\mathrm{H}_{2}^{+}$ entering the spiral inflector are as near to the central trajectory as possible. This is accomplished by the addition of electrostatic quadrupoles in the gap between the RFQ and inflector entrance (see also Fig. 8). Using the output of the RFQ simulation as input to the quadrupoles, geometry and potentials were optimized such that the tightest possible input beam is produced.
The remaining three modifications: V-shape, plate angling, and the terminal wedge were motivated by work done by Barnard et al. [60]. These all have the effect of introducing quadrupole moments to the field of the inflector itself, which can provide mid-track focusing to minimize beam dispersion in flight.
The V-shape introduces a concave v-shape to the anode and a convex v-shape to the cathode when viewed along the long axis of the electrode (see Fig. 33(a)). Plate angling reorients the plate angle relative to each other along the course of the inflector. In other words the plates are parallel only in the center with a positive angle between the plates initially linearly transitioning to an equal and opposite negative angle by the end (see Fig. 33(b)). The terminal gamma wedge cut modifies the end of the anode into a triangular rather than flat profile at the end and shortens the cathode. This assists with the fringe fields at the end which can produce vertical spreading. (see Fig. 33(c)).
@@
Figure 33: Illustrations of different focusing concepts. From Ref. [60].
@@
As we are at the edge of practical optimization using Monte Carlo, this leads us to consider other potential alternatives. Surrogate models using a fully connected neural network capable of learning and interpolating underlying features have been previously explored [61, 62, 63]. Initial iterations have proved promising and have informed the parameters for our current best inflector performance. As this work matures, it will be detailed in upcoming publications.